OpenStudy (anonymous):

line j: y = 1 4 x + 4 line k: y = 1 4 x + 3 line l: y = 3x + 4 line m: y = 4x + 1/4 These four lines have been graphed on the same coordinate grid. Which lines are parallel to each other?

jigglypuff314 (jigglypuff314):

Hello @Born_A_Genius16 and Welcome to OpenStudy! :) parallel means that the slopes would be equal to each other can you first identify what are the slopes of each given line? hint: when in slope intercept form y = mx + b "m" = slope

OpenStudy (anonymous):

line j: y = 1/4x + 4 line k: y = 1/4x +3 line l: y = 3x +4 line m: y = 4x +1/4